OpenStudy (anonymous):

Would f(x) = 4 sin (2x − π) - 1be 4cos(2x−(3π/2))-1 ??

OpenStudy (anonymous):

I was wondering how I could make this equation into a cos equation.

OpenStudy (anonymous):

@Kainui @bradely Do you guys understand what I'm saying?? :)

OpenStudy (bradely):

f(x) = 4 sin (2x − π) - 1 = 4 sin (-(π- 2x )) - 1 =-4sinsin (π- 2x ) - 1 =-4sin2x-1 f(x) =4cos(2x−(3π/2))-1 =4cos(-(-2x+(3π/2)))-1 =4cos(-2x+(3π/2))-1 =-4sin2x-1 So two functions are same Source: http://www.mathskey.com/question2answer/

OpenStudy (anonymous):

really?!?!? so would the vertex be -1?